OpenStudy (anonymous):

\[(1.3 x 10^{-6})(4 x 10^{10})\] Find the product and express your answer in scientific notation

sam (.sam.):

\[\Large (1.3 \times 10^{-6})(4\times 10^{10}) \\ \\ \Large =(1.3\times4)\times 10^{-6+10} \\ \\ \Large =5.2 \times 10^{4}\] Understand now?

OpenStudy (anonymous):

Indeed I do.

sam (.sam.):

okay :)

OpenStudy (anonymous):

Okay c: